OpenStudy (anonymous):

what property justifies the fallowing equation? 7x+5+2x=7x+2x+5

OpenStudy (anonymous):

Because 5+2x is the same as saying 2x+5, the equation is true.

OpenStudy (anonymous):

I meant .. Associative Property of Multiplication , Commutative Property of Addition , Commutative Property of Multiplication , Distributive Property , ?

OpenStudy (anonymous):

I believe it's Commutative property of Addition.

OpenStudy (nipunmalhotra93):

^yes it is

OpenStudy (anonymous):

thxs! i will be asking alot of questions I went from not caring about school to trying to do my best to graduate I hardly know this stuff :(

OpenStudy (anonymous):

To clear up why, let me give you some examples for the others: Associative property of multiplication: (2*3)*4=(2*3)*4 Commutative property of multiplication: 2*3=3*2 Distributive property: 2*(3*4)=(2*3)+(2*4)

OpenStudy (anonymous):

Also, you get a medal for doing your best. <3

OpenStudy (anonymous):

thnks thnks so much my lovelys!!